# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ IV.2.1 Z boson decays into a photon plus an unflavoured meson
The theoretical predictions and experimental limits for the branching fractions of the radiative Z boson decay into light, charm, and bottom unflavoured mesons ($\rm Z\to\gamma+\mathrm{q}\overline{\mathrm{q}},\mathrm{c}\overline{\mathrm{c}},\mathrm{b}\overline{\mathrm{b}}$, proceeding through the right and center diagrams of Fig. 15) are listed in Table 1112, and 13, respectively, and shown in graphical form in Fig. 16. After forty years from their discovery, no hadronic radiative decay of the electroweak bosons has yet been observed, despite searches performed by the CDF [139, 138], ATLAS [92, 140, 141, 99, 95], CMS [142, 143, 94, 96], and LHCb [144] experiments. In some cases, such as for the $\rm Z\to\gamma\eta,\,\gamma\eta^{\prime}$ channels, LEP measurements at the Z pole still provide the best upper limits [145].
The exclusive Z-boson radiative decays into light unflavoured mesons ($\rm Z\to\gamma+M$), listed in Table 11, have theoretical branching fractions in the range of $\mathcal{O}(10^{-8}$$10^{-12})$, whereas the current experimental limits are in the $\mathcal{O}(10^{-5}$$10^{-7})$ range with seven channels searched for by ATLAS (3), LHCb (1), CDF at Tevatron (1), and ALEPH at LEP (2). The $\rm Z\to\gamma\omega$ decay is very close to being detected at HL-LHC because its predicted $\mathcal{B}$ is about half of our projected conservative limit. Otherwise, all channels are producible at the FCC-ee.
@@
$\rm Z\,\to\,\gamma$ $+$ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee
$\rm Z\,\to\,\gamma$ + $\rm\pi^{0}$  $(9.8\pm 1.0)$$\times~{}10^{-12}$ SCET+LCDA [44] $<$ $2.0$×$\times~{}10^{-5}$ [138] $\lesssim 2.8\times~{}10^{-7}$
@@
$\rm\rho^{0}$  $(4.19\pm 0.47)$$\times~{}10^{-9}$ SCET+LCDA [44] $<$ $4.0$×$\times~{}10^{-6}$ [92] $\lesssim 2.9\times~{}10^{-7}$
$\rm\omega$  $(2.82\pm 0.41)$$\times~{}10^{-8}$ SCET+LCDA [44] $<$ $3.9$×$\times~{}10^{-6}$ [95] $\lesssim 5.8\times~{}10^{-7}$
@@
$\rm\phi$  $(1.04\pm 0.12)$$\times~{}10^{-8}$ SCET+LCDA [44] $<$ $7.0$×$\times~{}10^{-7}$ [92] $\lesssim 5.1\times~{}10^{-8}$
 $(1.17\pm 0.08)$$\times~{}10^{-8}$ LC+LCDA [147]
The exclusive Z-boson radiative decays to charmonium mesons ($\rm Z\to\gamma+\mathrm{c}\overline{\mathrm{c}}$), collected in Table 12, have rates in the $\mathcal{O}(10^{-8}$$10^{-10})$ range as computed within multiple approaches (LC, SCET, NRQCD), which are in general consistent among themselves for the same process. Only one single channel has been searched for, $\rm Z\to\gamma+\mathrm{J}/\psi$, with $\mathcal{O}(10^{-6})$ experimental upper bounds [96, 142]. This decay may be visible at HL-LHC, because its predicted rate is about $1/5$ of our projected conservative limit. All channels are producible at FCC-ee.
@@
$\rm Z\,\to\,\gamma$ $+$ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee
$\rm Z\,\to\,\gamma$ + $\rm\eta_{c}$  $\left(1.32^{+0.56}_{-0.54}\right)$$\times~{}10^{-8}$ NRQCD+LDME [148]
 $(9.5\pm 0.2)$$\times~{}10^{-9}$ NRQCD (NNLO+NLL) [149]
 $(7.42\pm 0.61)$$\times~{}10^{-9}$ NRQCD (NLO+NLL) [150]
 $6.6$$\times~{}10^{-9}$ NRQCD+LDME [151]
 $(9.4\pm 1.0)$$\times~{}10^{-9}$ LC+LCDA [151]
$\rm J/\psi$  $\left(1.04^{+0.18}_{-0.16}\right)$$\times~{}10^{-7}$ NRQCD+LDME [148] $<$ $6.0$×$\times~{}10^{-7}$ [96, 142] $\lesssim 3.1\times~{}10^{-7}$ [60]
 $\left(5.75^{+0.08}_{-0.09}\right)$$\times~{}10^{-8}$ NRQCD (NNLO+NLL) [149]
 $\left(8.96^{+1.51}_{-1.38}\right)$$\times~{}10^{-8}$ LC+LCDA [152]
 $4.5$$\times~{}10^{-8}$ NRQCD+LDME [151]
 $(8.8\pm 0.9)$$\times~{}10^{-8}$ LC+LCDA [151]
 $(8.02\pm 0.45)$$\times~{}10^{-8}$ SCET+LCDA [44]
 $(9.96\pm 1.86)$$\times~{}10^{-8}$ NRQCD+LDME [147]
$\rm\psi(2S)$  $\left(4.83^{+1.02}_{-0.91}\right)$$\times~{}10^{-8}$ SCET+LCDA999private communication with author of Ref. [100] [96] $<$ $1.3$×$\times~{}10^{-6}$ [96, 142] $\lesssim 1.0\times~{}10^{-7}$
$\rm\chi_{c0}$  $(3.74\pm 0.05)$$\times~{}10^{-10}$ NRQCD+LDME [153]
 $1.4$$\times~{}10^{-10}$ NRQCD+LDME [151]
 $(5.0\pm 2.0)$$\times~{}10^{-10}$ LC+LCDA [151]
$\rm\chi_{c1}$  $\left(2.383^{+0.014}_{-0.017}\right)$$\times~{}10^{-9}$ NRQCD+LDME [153]
 $8.7$$\times~{}10^{-10}$ NRQCD+LDME [151]
 $(5.6\pm 2.0)$$\times~{}10^{-9}$ LC+LCDA [151]
$\rm h_{c}$  $\left(3.487^{+0.206}_{-0.230}\right)$$\times~{}10^{-9}$ NRQCD+LDME [153]
 $3.0$$\times~{}10^{-9}$ NRQCD+LDME [151]
 $(1.0\pm 0.4)$$\times~{}10^{-8}$ LC+LCDA [151]
$\rm\chi_{c2}$  $\left(3.38^{+0.19}_{-0.22}\right)$$\times~{}10^{-10}$ NRQCD+LDME [153]
 $2.9$$\times~{}10^{-10}$ NRQCD+LDME [151]
 $(1.0\pm 0.4)$$\times~{}10^{-9}$ LC+LCDA [151]
@@
$\rm Z\,\to\,\gamma$ $+$ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee
$\rm Z\,\to\,\gamma$ + $\rm\eta_{b}$  $\left(2.79^{+0.20}_{-0.19}\right)$$\times~{}10^{-8}$ NRQCD+LDME [148]
 $(2.43\pm 0.01)$$\times~{}10^{-8}$ NRQCD (NNLO+NLL) [149]
 $(2.8\pm 0.5)$$\times~{}10^{-8}$ NRQCD (NLO+NLL) [150]
$\rm\Upsilon(1S)$  $(5.41\pm 0.39)$$\times~{}10^{-8}$ NRQCD+LDME [148] $<$ $1.1$×$\times~{}10^{-6}$ [99] $\lesssim 1.7\times~{}10^{-7}$
 $(4.63\pm 0.02)$$\times~{}10^{-8}$ NRQCD (NNLO+NLL) [149]
 $(5.61\pm 0.29)$$\times~{}10^{-8}$ NRQCD+LDME [154]
 $\left(4.8^{+0.3}_{-0.2}\right)$$\times~{}10^{-8}$ LC+LCDA [152]
 $(5.39\pm 0.16)$$\times~{}10^{-8}$ SCET+LCDA [44]
 $(4.93\pm 0.51)$$\times~{}10^{-8}$ NRQCD+LDME [147]
$\rm\Upsilon(2S)$  $(2.66\pm 0.31)$$\times~{}10^{-8}$ NRQCD+LDME [154] $<$ $1.3$×$\times~{}10^{-6}$ [99] $\lesssim 2.0\times~{}10^{-7}$
 $\left(2.44^{+0.14}_{-0.13}\right)$$\times~{}10^{-8}$ LC+LCDA [152]
$\rm\Upsilon(3S)$  $(1.93\pm 0.25)$$\times~{}10^{-8}$ NRQCD+LDME [154] $<$ $2.4$×$\times~{}10^{-6}$ [99] $\lesssim 3.7\times~{}10^{-7}$
 $\left(1.88^{+0.11}_{-0.10}\right)$$\times~{}10^{-8}$ LC+LCDA [152]
$\rm\Upsilon(4S)$  $(1.22\pm 0.13)$$\times~{}10^{-8}$ SCET+LCDA [44]
$\rm\Upsilon(nS)$  $\left(9.96^{+0.28}_{-0.26}\right)$$\times~{}10^{-8}$ SCET+LCDA [44]
$\rm\chi_{b0}$  $\left(2.73^{+0.05}_{-0.04}\right)$$\times~{}10^{-10}$ NRQCD+LDME [153]
$\rm\chi_{b1}$  $\left(1.473^{+0.010}_{-0.011}\right)$$\times~{}10^{-9}$ NRQCD+LDME [153]
$\rm h_{b}$  $\left(9.27^{+0.36}_{-0.41}\right)$$\times~{}10^{-10}$ NRQCD+LDME [153]
$\rm\chi_{b2}$  $\left(2.92^{+0.12}_{-0.14}\right)$$\times~{}10^{-10}$ NRQCD+LDME [153]
Figure 16: Branching fractions (in negative log scale) of exclusive $\rm Z\to\gamma+M$ decays, with M being light (upper), charm (middle), and bottom (lower) unflavoured mesons. The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-ee reach based only on the total number of Z bosons to be produced.
@@ IV.2.2 Z boson decays into a photon plus a flavoured meson
In the SM, the absence of direct FCNC decays of the Z boson implies that the only possible exclusive radiative Z boson decays into flavoured mesons can proceed through the (extremely suppressed) right diagram of Fig. 15. We are not aware of existing evaluations of the branching fractions for such exclusive decays in the SM, but only of indirect limits for them determined in Ref. [44] using different SM constraints. Therefore, we estimate the numerical value of these exclusive FCNC channels here next. The standard Lagrangian density describing the Z boson coupling to quarks is given by the interaction term
@@
where $i,j\in\{\rm u,d,c,s,t,b\}$. The offdiagonal interaction couplings $v_{\overline{i},j\neq i}$ and $a_{\overline{i},j\neq i}$ are zero in the SM, but are used here as effective couplings to encode contributions from higher-order loop processes (such as those shown in Fig. 15, right), which can take on a very small value. The work [44] provides generic expressions for the calculation of partial widths of exclusive Z boson decays into a flavoured meson plus a photon in the SCET+\,+\,+LCDA framework, using offdiagonal quark couplings $v_{\overline{i}j},a_{\overline{i}j}$ as inputs. For arbitrary $\rm\overline{q_{i}}q_{j}$ combinations, the decay width for the $\rm Z\to\overline{q}q^{\prime}$ process can be written at tree level as
@@
Using this last expression, the numerical values of the effective couplings $v_{\mathrm{\overline{q}q^{\prime}}}=a_{\mathrm{\overline{q}q^{\prime}}}$ can now be estimated from the offdiagonal Z boson quark decay widths computed at NLO accuracy in [91]: $\Gamma_{\mathrm{Z\to\bar{b}s}}=5.2\times 10^{-8}$ GeV, $\Gamma_{\mathrm{Z\to\bar{b}d}}=2.3\times 10^{-9}$ GeV, $\Gamma_{\mathrm{Z\to\bar{c}u}}=1.7\times 10^{-18}$ GeV; and amount to
@@
The decay $\rm Z\to\overline{s}d$ was first computed in [155], with vector and axial couplings equal since the photon contribution is zero, and with the offdiagonal coupling obtained via
@@
Plugging in the values of the effective couplings (38) and (40) into the SCET+\,+\,+LCDA expressions of Ref. [44] enables us to determine the numerical branching fractions for Z-boson radiative decays to flavoured mesons listed in Table 14 and graphically presented in Fig. 17. The quoted rates are for the sum of $\rm M=M^{0}+\overline{M^{0}}$ radiative decays. All such exclusive FCNC decays of the Z boson are extremely suppressed, amounting to $\mathcal{O}(10^{-15}\mbox{--}10^{-25})$. The experimentally negligible branching fractions for such decays make those interesting channels to search for BSM FCNC decays of the Z boson without any irreducible SM background. Two decays have been experimentally searched-for to date: $\rm Z\to\gamma+D^{0}$ [144, 115] and $\rm Z\to\gamma+K_{s}^{0}$ [115], where experimental limits are at the $\mathcal{O}(10^{-6})$ level, and can be improved by a factor of $\sim$10 at the end of the HL-LHC phase as per our extrapolations here.
@@
$\rm Z\,\to\,\gamma$ $+$ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee
$\rm Z\,\to\,\gamma$ + $\rm K^{0}$  $3.3$$\times~{}10^{-20}$ SCET+LCDA (this work) $<$ $3.0$×$\times~{}10^{-6}$ [115] $\lesssim 4.5\times~{}10^{-7}$
$\rm D^{0}$ $1.0$$\times~{}10^{-15}$ SCET+LCDA [44, 144] $<$ $4.0$×$\times~{}10^{-6}$ [115] $\lesssim 6.0\times~{}10^{-7}$
@@
Figure 17: Branching fractions (in negative log scale) of exclusive $\rm Z\to\gamma+M$ decays, with M being a flavoured meson. The theoretical predictions computed here (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-ee reach based only on the total number of Z bosons to be produced.

### v4

@@ IV.2.1 Z boson decays into a photon plus an unflavoured meson
The theoretical predictions and experimental limits for the branching fractions of the radiative Z boson decay into light, charm, and bottom unflavoured mesons ($\rm Z\to\gamma+\mathrm{q}\overline{\mathrm{q}},\mathrm{c}\overline{\mathrm{c}},\mathrm{b}\overline{\mathrm{b}}$, proceeding through the left and center diagrams of Fig. 15) are listed in Table 1112, and 13, respectively, and shown in graphical form in Fig. 16. After forty years from their discovery, no hadronic radiative decay of the electroweak bosons has yet been observed, despite searches performed by the CDF [CDF:1998kzn, CDF:2013lma], ATLAS [ATLAS:2017gko, ATLAS:2018jqi, ATLAS:2018zsq, ATLAS:2022rej, ATLAS:2023alf], CMS [CMS:2018gcm, CMS:2019vaj, CMS:2019wch, CMS:2024hhg], and LHCb [LHCb:2022kta] experiments. In some cases, such as for the $\rm Z\to\gamma\eta,\,\gamma\eta^{\prime}$ channels, LEP measurements at the Z pole still provide the best upper limits [ALEPH:1991qhf].
The exclusive Z-boson radiative decays into light unflavoured mesons ($\rm Z\to\gamma+M$), listed in Table 11, have theoretical branching fractions in the range of $\mathcal{O}(10^{-8}$$10^{-12})$, whereas the current experimental limits are in the $\mathcal{O}(10^{-5}$$10^{-7})$ range with seven channels searched for by ATLAS (3), LHCb (1), CDF at Tevatron (1), and ALEPH at LEP (2). The $\rm Z\to\gamma\phi$ decay is close to being detected at HL-LHC because its predicted $\mathcal{B}$ is about 1/5 of our projected conservative limit. Otherwise, all channels are producible at the FCC-ee.
@@
$\rm Z\,\to\,\gamma$ ++ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$
@@
[…]
$\rm\phi$ $(1.04\pm 0.12)$ $\times~10^{-8}$ SCET+LCDA [Grossman:2015cak] << $07\times~10^{-7}$ [ATLAS:2017gko] $\lesssim 5.1\times~10^{-8}$
$(1.17\pm 0.08)$ $\times~10^{-8}$ LC+LCDA [Huang:2014cxa]
The exclusive Z-boson radiative decays to charmonium mesons ($\rm Z\to\gamma+\mathrm{c}\overline{\mathrm{c}}$), collected in Table 12, have rates in the $\mathcal{O}(10^{-8}$$10^{-10})$ range as computed within multiple approaches (LC, SCET, NRQCD), which are in general consistent among themselves for the same process. Only one single channel has been searched for, $\rm Z\to\gamma+\mathrm{J}/\psi$, with $\mathcal{O}(10^{-6})$ experimental upper bounds [CMS:2024hhg, CMS:2018gcm]. This decay may be visible at HL-LHC, because its predicted rate is about $1/3$ of our projected conservative limit. All channels are producible at FCC-ee.
@@
$\rm Z\,\to\,\gamma$ ++ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$
$\rm Z\,\to\,\gamma$ + $\rm\eta_{c}$ $\left(1.32^{+0.56}_{-0.54}\right)$ $\times~10^{-8}$ NRQCD+LDME [Wang:2023ssg]
$(9.5\pm 0.2)$ $\times~10^{-9}$ NRQCD (NNLO+NLL) [Sang:2023hjl]
$(7.42\pm 0.61)$ $\times~10^{-9}$ NRQCD (NLO+NLL) [Chung:2019ota]
$6.6$ $\times~10^{-9}$ NRQCD+LDME [Luchinsky:2017jab]
$(9.4\pm 1.0)$ $\times~10^{-9}$ LC+LCDA [Luchinsky:2017jab]
[…]
$\left(5.75^{+0.08}_{-0.09}\right)$ $\times~10^{-8}$ NRQCD (NNLO+NLL) [Sang:2023hjl]
$\left(8.96^{+1.51}_{-1.38}\right)$ $\times~10^{-8}$ LC+LCDA [Bodwin:2017pzj]
$4.5$ $\times~10^{-8}$ NRQCD+LDME [Luchinsky:2017jab]
$(8.8\pm 0.9)$ $\times~10^{-8}$ LC+LCDA [Luchinsky:2017jab]
$(8.02\pm 0.45)$ $\times~10^{-8}$ SCET+LCDA [Grossman:2015cak]
$(9.96\pm 1.86)$ $\times~10^{-8}$ NRQCD+LDME [Huang:2014cxa]
[…]
$\rm\chi_{c0}$ $(3.74\pm 0.05)$ $\times~10^{-10}$ NRQCD+LDME [Sang:2022erv]
$1.4$ $\times~10^{-10}$ NRQCD+LDME [Luchinsky:2017jab]
$(5.0\pm 2.0)$ $\times~10^{-10}$ LC+LCDA [Luchinsky:2017jab]
$\rm\chi_{c1}$ $\left(2.383^{+0.014}_{-0.017}\right)$ $\times~10^{-9}$ NRQCD+LDME [Sang:2022erv]
$8.7$ $\times~10^{-10}$ NRQCD+LDME [Luchinsky:2017jab]
$(5.6\pm 2.0)$ $\times~10^{-9}$ LC+LCDA [Luchinsky:2017jab]
$\rm h_{c}$ $\left(3.487^{+0.206}_{-0.230}\right)$ $\times~10^{-9}$ NRQCD+LDME [Sang:2022erv]
$3.0$ $\times~10^{-9}$ NRQCD+LDME [Luchinsky:2017jab]
$(1.0\pm 0.4)$ $\times~10^{-8}$ LC+LCDA [Luchinsky:2017jab]
$\rm\chi_{c2}$ $\left(3.38^{+0.19}_{-0.22}\right)$ $\times~10^{-10}$ NRQCD+LDME [Sang:2022erv]
$2.9$ $\times~10^{-10}$ NRQCD+LDME [Luchinsky:2017jab]
$(1.0\pm 0.4)$ $\times~10^{-9}$ LC+LCDA [Luchinsky:2017jab]
@@
$\rm Z\,\to\,\gamma$ ++ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$
$\rm Z\,\to\,\gamma$ + $\rm\eta_{b}$ $\left(2.79^{+0.20}_{-0.19}\right)$ $\times~10^{-8}$ NRQCD+LDME [Wang:2023ssg]
$(2.43\pm 0.01)$ $\times~10^{-8}$ NRQCD (NNLO+NLL) [Sang:2023hjl]
$(2.8\pm 0.5)$ $\times~10^{-8}$ NRQCD (NLO+NLL) [Chung:2019ota]
[…]
$(4.63\pm 0.02)$ $\times~10^{-8}$ NRQCD (NNLO+NLL) [Sang:2023hjl]
$(5.61\pm 0.29)$ $\times~10^{-8}$ NRQCD+LDME [Dong:2022ayy]
$\left(4.8^{+0.3}_{-0.2}\right)$ $\times~10^{-8}$ LC+LCDA [Bodwin:2017pzj]
$(5.39\pm 0.16)$ $\times~10^{-8}$ SCET+LCDA [Grossman:2015cak]
$(4.93\pm 0.51)$ $\times~10^{-8}$ NRQCD+LDME [Huang:2014cxa]
[…]
$\left(2.44^{+0.14}_{-0.13}\right)$ $\times~10^{-8}$ LC+LCDA [Bodwin:2017pzj]
[…]
$\left(1.88^{+0.11}_{-0.10}\right)$ $\times~10^{-8}$ LC+LCDA [Bodwin:2017pzj]
$\rm\Upsilon(4S)$ $(1.22\pm 0.13)$ $\times~10^{-8}$ SCET+LCDA [Grossman:2015cak]
$\rm\Upsilon(nS)$ $\left(9.96^{+0.28}_{-0.26}\right)$ $\times~10^{-8}$ SCET+LCDA [Grossman:2015cak]
$\rm\chi_{b0}$ $\left(2.73^{+0.05}_{-0.04}\right)$ $\times~10^{-10}$ NRQCD+LDME [Sang:2022erv]
$\rm\chi_{b1}$ $\left(1.473^{+0.010}_{-0.011}\right)$ $\times~10^{-9}$ NRQCD+LDME [Sang:2022erv]
$\rm h_{b}$ $\left(9.27^{+0.36}_{-0.41}\right)$ $\times~10^{-10}$ NRQCD+LDME [Sang:2022erv]
$\rm\chi_{b2}$ $\left(2.92^{+0.12}_{-0.14}\right)$ $\times~10^{-10}$ NRQCD+LDME [Sang:2022erv]
Figure 16: Branching fractions (in negative log scale) of exclusive $\rm Z\to\gamma+M$ decays, with M being light (upper), charm (middle), and bottom (lower) unflavoured mesons. The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-$e$ reach based only on the total number of Z bosons to be produced.
@@ IV.2.2 Z boson decays into a photon plus a flavoured meson
In the SM, the absence of direct FCNC decays of the Z boson implies that the only possible exclusive radiative Z boson decays into flavoured mesons can proceed through the (extremely suppressed) right diagram of Fig. 15. We are not aware of existing evaluations of the branching fractions for such exclusive decays in the SM, but only of indirect limits for them determined in Ref. [Grossman:2015cak] using different SM constraints. Therefore, we estimate the numerical value of these exclusive FCNC channels here next. The standard Lagrangian density describing the Z boson coupling to quarks is given by the interaction term
@@
where $i,j\in\{\rm u,d,c,s,t,b\}$. The offdiagonal interaction couplings $v_{\overline{i},j\neq i}$ and $a_{\overline{i},j\neq i}$ are zero in the SM, but are used here as effective couplings to encode contributions from higher-order loop processes (such as those shown in Fig. 15, right), which can take on a very small value. The work [Grossman:2015cak] provides generic expressions for the calculation of partial widths of exclusive Z boson decays into a flavoured meson plus a photon in the SCET+\,+\,LCDA framework, using offdiagonal quark couplings $v_{\overline{i}j},a_{\overline{i}j}$ as inputs. For arbitrary $\rm\overline{q_{i}}q_{j}$ combinations, the decay width for the $\rm Z\to\overline{q}q^{\prime}$ process can be written at tree level as
@@
Using this last expression, the numerical values of the effective couplings $v_{\mathrm{\overline{q}q^{\prime}}}=a_{\mathrm{\overline{q}q^{\prime}}}$ can now be estimated from the offdiagonal Z boson quark decay widths computed at NLO accuracy in [Kamenik:2023hvi]: $\Gamma_{\mathrm{Z\to\bar{b}s}}=5.2\times 10^{-8}$ GeV, $\Gamma_{\mathrm{Z\to\bar{b}d}}=2.3\times 10^{-9}$ GeV, $\Gamma_{\mathrm{Z\to\bar{c}u}}=1.7\times 10^{-18}$ GeV; and amount to
@@
The decay $\rm Z\to\overline{s}d$ was first computed in [Ma:1979px], with vector and axial couplings equal since the photon contribution is zero, and with the offdiagonal coupling obtained via
@@
Plugging in the values of the effective couplings (38) and (40) into the SCET+\,+\,LCDA expressions of Ref. [Grossman:2015cak] enables us to determine the numerical branching fractions for Z-boson radiative decays to flavoured mesons listed in Table 14 and graphically presented in Fig. 17. The quoted rates are for the sum of $\rm M=M^{0}+\overline{M^{0}}$ radiative decays. All such exclusive FCNC decays of the Z boson are extremely suppressed, amounting to $\mathcal{O}(10^{-15}\mbox{--}10^{-25})$. The experimentally negligible branching fractions for such decays make those interesting channels to search for BSM FCNC decays of the Z boson without any irreducible SM background. Two decays have been experimentally searched-for to date: $\rm Z\to\gamma+D^{0}$ [LHCb:2022kta, ATLAS:2024dpw] and $\rm Z\to\gamma+K_{s}^{0}$ [ATLAS:2024dpw], where experimental limits are at the $\mathcal{O}(10^{-6})$ level, and can be improved by a factor of $\sim$10 at the end of the HL-LHC phase as per our extrapolations here.
@@
$\rm Z\,\to\,\gamma$ ++ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$
[…]
@@
Figure 17: Branching fractions (in negative log scale) of exclusive $\rm Z\to\gamma+M$ decays, with M being a flavoured meson. The theoretical predictions computed here (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-$e$ reach based only on the total number of Z bosons to be produced.